x + 3( x + 1 )= 2x
Answer by pwac(253)

x+3(x+1)=2x
First remove bracket by multiplying by the 3
x+3x+3=2x Subtract 3 from both sides
x+3x=2x-3 Collect x terms on L.H.S
x+3x-2x=-3 (remember change the side change the sign.Sum x terms
2x=-3 divide both sides by 2gives
x=-3/2 or -1.5
Substitute into original equation to check.
It works -3=-3